# From Observation to Orientation: an Adaptive Integer Programming Approach to Intervention Design

### v1

@@ 4.3.1 Transition Rules for Edge Sets
If the edge $(i,j)$ is subject to an orientation test ($O_{ij}=1$ or $O_{ji}=1$) and the intervention reveals the edge to be present, it is moved to $E_{\text{Known}}$ with its resolved direction. The edge is then removed from $E_{\text{Adjacent}}$.
5 Theoretical Analysis
@@
In this section, we provide theoretical guarantees for the proposed adaptive intervention design algorithm (Algorithm 1). These results hold under the standard assumptions commonly made in causal discovery from interventions:
@@
6 Experiments
To evaluate the performance of the Adaptive_IP model, we conduct a series of simulations using synthetic data. The primary goal of these experiments is to assess the model’s ability to recover the true causal structure of a DAG under different conditions, including varying graph sizes, edge densities, and intervention limits. For each simulation, the initial PKG is the essential graph of the true graph.
6.1 Experimental Setup
6.1.1 Data Generation
For each simulation, we generate a random DAG using the $\mathcal{G}(N,p)$ Erdős-Rényi model. Specifically, this model generates an undirected graph with $N$ nodes, where each edge is created independently with a probability $p$. The edges are then directed according to the identity topological ordering of the nodes to ensure acyclicity.
6.1.2 Simulation Parameters
The following parameters are systematically varied across the simulations:
  • Number of Nodes ($N$): Graphs with $N\in\{3,4,8,16,24,32,48,64,96,128,256\}$ are experimented.
  • Edge Probability ($p$): Edge probabilities $p$ are chosen from a set of $\{0.05,0.2,0.5,0.7,0.95\}$ to control graph density.
  • Maximum Interventions ($k_{max}$): We limited the maximum number of simultaneous interventions in each iteration to $k_{max}\in\{1,2,4,6\}$. This allows us to study the impact of intervention capacity on causal discovery.
  • Methods: The performances of two methods are compared:
    • ‘IP’: Adaptive_IP
6.1.3 Metrics
The following metrics are recorded from each experiment:
  • Total Runtime: The total runtime of the Adaptive_IP model.
Simulations are implemented in Python using the NetworkX library for graph generation and the Gurobi solver for solving the Adaptive_IP model. Each scenario is repeated for 50 random graphs. For the purpose of this study, we assume that the cost of interventions is negligible and hence constraint 2 is not added to the model. Additionally, for instances where multiple optimal solutions to the Adaptive_IP model exist, a random one is chosen.
6.2 Results and Discussion
The results of the simulations are presented in Figure 2, showing the mean number of iterations required for convergence as a function of graph density ($p$) for different values of $N$ and $k_{max}$. The results are grouped by the two methods.
Figure 2: Mean number of iterations required for convergence as a function of graph density ($p$) for different values of $N$, and $k_{max}$. The left column shows the results for the ‘r’ method and the right row shows the results for the ‘IP’ method.
Figure 3: a) Scatterplot of the number of iterations vs. the number of v-structures/$\binom{N}{3}$ by $k_{max}$. b) Mean number of v-structures vs. $p$ by $N$.
A key observation from Figure 2 is that the ‘IP’ method consistently outperforms the ‘r’ method, requiring fewer iterations to learn the graph across all values of $N$, $p$, and $k_{max}$ on average. This suggests that the proposed approach is more efficient in the context of structure learning. Furthermore, the error bars, which represent the range between the minimum and maximum number of iterations, are generally larger for the ‘r’ method. The difference in number of iterations required by both methods becomes smaller as $k_{max}$ becomes larger. However, the ‘IP’ method required fewer variable manipulations to learn the causal structure as shown in Figure 4 in Appendix.
Another crucial observation is the non-monotonic relationship between the number of iterations and graph density for all values of $k_{max}$. This relationship is particularly pronounced for lower $k_{max}$ values. Starting from very low densities ($p=0.05$), the number of iterations initially decreases as density increases, reaching a minimum before increasing again as the density continues to rise. This trend can be explained by considering the interplay between v-structures, maximal cliques, and the application of Meek’s rules during the learning process. Specifically, at very low densities, the true graph is sparse, and consequently, there are very few v-structures/maximal cliques. Thus, the algorithm has limited opportunities to apply Meek’s rules for edge orientation. As the density slightly increases, the number of v-structures and maximal cliques begins to grow. Interventions can then reveal these substructures, enabling the application of Meek’s rules to orient more edges. This leads to a decrease in the number of iterations required compared to the very sparse case. However, as the density increases further, the number of maximal cliques in the graph continues to increase. This shields the colliders involved in the v-structures, which decreases the ability of Meek’s rules to orient edges, thus more interventions are needed to resolve the orientations. These justifications are supported by Figure 3.
For larger values of $k_{max}$, the algorithm can resolve most of the orientations by only a few iterations. More interestingly, the number of interventions required to learn a graph starting from its essential graph is constant in expectation (between 1 and 3 iterations) and independent of the graph’s number of nodes/edges. This result aligns with the theoretical bound formally shown in [24].
7 Conclusion
In this paper, we proposed an adaptive approach to the design of intervention experiments for causal discovery. Our approach uses an IP approach to select the most informative interventions given a set of practical constraints such as a bound on the total budget and the number of simultaneous interventions sequentially. We further reduce the experimental burden by incorporating Meek’s rules, which propagate additional orientations without extra interventions. In future work, this framework can be extended to operate in a noisy data setting as well as settings with unmeasured confounders. This work underlines the potential of adaptive cost-efficient strategies for improving causal discovery.
Appendix A Appendix
Figure 4: Mean number of total variables maniupulations for convergence as a function of graph density ($p$) for different values of $N$, and $k_{max}$. The left column shows the results for the ‘r’ method and the right row shows the results for the ‘IP’ method.
Figure 5: Run time (s) as a function of graph density ($p$) for different values of $N$ and $k_{max}$.

### v3

@@ 3.2 Transition Rules for Edge Sets
If the edge $(i,j)$ is subject to an orientation test and the intervention reveals the edge to be present, it is moved to $E_{\text{Known}}$. The edge is then removed from $E_{\text{Adjacent}}$.
4 Experiments
To evaluate the performance of Algorithm 1, we conduct a series of simulations using synthetic graphs as well as real graphs. The primary goal of these experiments is to assess the model’s ability to learn the true DAG under different conditions such as graph sizes, edge densities, and intervention size limits. While the algorithm can work for any initial PKG, for the purpose of this study, the initial PKG is made to be the correct essential graph of the true graph.
4.1 Synthetic Graphs
For each simulation, we generate 50 random DAGs using the $\mathcal{G}(N,p)$ Erdős-Rényi model. This model generates an undirected graph with $N$ nodes, where each edge is created independently with a probability $p$. The edges are then directed according to the identity topological ordering of the nodes to ensure acyclicity. This is done for the set of parameters: $\{(N,p)\ |\ N\in\{3,4,8,16,24,32,48,64,96,128,256\},\ p\in\{0.05,0.2,0.5,0.7,0.95\}\}$.
4.2 Real Graphs
Algorithm 1 was tested against 9 real graphs. The graphs have different sizes, densities, and degree distributions. Table 2 summarizes the structural statistics of each graph.
Table 2: Structural statistics of benchmark Graphs
Network Nodes Edges Min deg. Avg. deg. Max deg. Stdev. deg. # v-structures
asia 8 8 1 2.00 4 0.93 2
sachs 11 17 2 3.09 7 1.64 0
insurance 27 52 1 3.85 9 2.03 23
alarm 37 46 1 2.49 6 1.35 24
hailfinder 56 66 1 2.36 17 2.40 34
win95pts 76 112 1 2.95 10 2.01 129
pathfinder 109 195 1 3.58 106 10.16 16
andes 223 338 0 3.03 12 1.87 313
link 724 1125 0 3.11 17 2.49 821
4.2.1 Comparison Parameters and Metrics
The following parameters are varied across the simulations:
  • Maximum Interventions ($k_{max}$): We limited the maximum number of simultaneous interventions in each iteration to $k_{max}\in\{1,2,4,6\}$.
  • Methods: The performances of two methods are compared:
    • ‘IP’: Adaptive_IP
The following metrics are recorded from each experiment:
For the purpose of this study, we assume that the cost of interventions is negligible and hence constraint 2 is not added to the model. Additionally, for instances where multiple optimal solutions to the Adaptive_IP model exist, a random one is chosen. Simulations are implemented in Python using the NetworkX [31] library for synthetic graph generation and the Gurobi solver for solving the Adaptive_IP model.
4.3 Results and Discussion
Figures 2 and 3 plot, for every experimental setting, the quantity
$\Delta_{rounds}\;=\;(\#\text{rounds}_{\textit{r}})-(\#\text{rounds}_{\textit{Adaptive\_IP}})$
that is the additional intervention iterations required by the random baseline. Hence positive values indicate an unambiguous advantage for the optimized strategy, while $\Delta_{rounds}=0$ corresponds to parity. Across all experiments median $\Delta_{rounds}\geq 0$, confirming that Adaptive_IP never performs worse than random selection and often yields substantial savings.
Synthetic Erdős-Rényi graphs
(Figure 2). The magnitude of the gap depends strongly on graph size $N$, edge probability $p$, and the per-round intervention budget $k_{\max}$:
$k_{max}$ dependence For each facet, when $k_{\max}=6$, the distributions sit highest for ($0.20\leq p\leq 0.70$): when many variables can be perturbed in parallel, Adaptive_IP can exploit the large combinatorial space to pick a particularly informative six-tuple, whereas random choice often squanders part of that budget. In contrast, for the dense regime ($p=0.95$) the benefit is largest when only one variable can be manipulated. There the baseline’s single random knock-out is frequently uninformative, whereas the IP solver consistently targets a key hub; as $k_{\max}$ increases the gap collapses because almost any two or more perturbations break enough cycles to let Meek’s rules finish the job.
Density dependence Very sparse graphs ($p=0.05$) and very dense graphs ($p=0.95$) exhibit the broadest inter-quartile ranges—up to twelve rounds in the sparse case and more than sixty rounds in the dense case—because each contains few v-structures, leaving Meek’s propagation with little to do; progress hinges on how informative the chosen interventions are. Intermediate densities ($0.20\leq p\leq 0.70$) feature many v-structures, hence once a handful of orientations are fixed Meek’s rules cascade through the graph and even randomly chosen interventions achieve respectable coverage; the median advantage in these panels rarely exceeds ten rounds.
Figure 2: Number of additional intervention rounds required when using method $r$ versus Adaptive_IP. Each panel corresponds to a graph-edge probability $p$; bars show the distribution of $r-\mathrm{IP}$ across seeds for different intervention limits $k_{\max}$ and network sizes $N$.
Real benchmark graphs.
Figure 3 compares the extra intervention rounds required by the random baseline across the nine real networks whose structural statistics are listed in Table 2. interventions are allowed. The extent of the advantage of the adaptive IP framework differs sharply from graph to graph and tracks well with the topological indicators in Table 2. In pathfinder the median saving exceeds forty iterations; the network contains a central hub node (maximum degree 106) and progress hinges on perturbing that specific node—something the IP identifies reliably but a random pick rarely does. hailfinder and insurance also benefit markedly ($\Delta_{rounds}\approx 4\!-\!8$) because their degree distributions are skewed and they possess only a moderate number of v-structures (34 and 23, respectively), leaving Meek’s rules with limited reach unless the right edges are experimentally oriented. At the other extreme, link shows virtually no median improvement even at $k_{\max}=1$; with 821 v-structures, many edges are already oriented in the MEC, and the random baseline performs as well as the optimized selection. Small sparse graphs such as asia and sachs exhibit the same phenomenon on a lesser scale—the uncertainty is so limited that even a naïve intervention suffices to complete the discovery in just one or two additional rounds. Taken together, the real-graph results highlight when an adaptive IP planner is worth its computational cost: namely, in networks whose orientation hinges on a few critical hubs or on edges that cannot be resolved by Meek’s rules alone, and in experimental regimes where only one variable can be perturbed at a time. When many v-structures are present—or when even a modest multi-variable budget is available—the simpler random strategy approaches parity.
Figures 4 and 5 demonstrate similar results for the number of additional number of variable-manipulations required for graph identification by the random baseline.
These simulations aim to validate the feasibility of Algorithm 1 under the assumptions listed in Section 3. While the focus of the simulations conducted was on the efficiency of the selected interventions in learnine the true graph, we emphasize that the reported benefit of the framework is attributed to the nature of Objective 1. There exists other objectives (see Section C.2) that can outperform Objective 1 (which is myopic), albeit, tend to be more computationally expensive to evaluate. Furthermore, we emphasize that the main contribution of this work is its offered modularity and flexibility. For instance, the same framework can be utilized with an alternative objective while accommodating practical constraints such as intervention costs and experimentation budget, limits on the number of variables that can be intervened on per iteration, prior knowledge, and other considerations by simply replacing Objective 1 with another. No other adaptive intervention design strategy offers this level of flexibility and hence it was difficult to directly compare the proposed framework to methods found in the literature.
5 Conclusion
This paper presented a novel adaptive intervention strategy for causal discovery built upon iterative Integer Programming. A primary contribution is the IP framework’s flexibility, allowing seamless integration of practical constraints and diverse objectives, which is challenging for existing methods. Simulations demonstrated performance improvements over random baselines, and theoretical analysis confirmed finite convergence and correctness under assumptions including Causal Sufficiency and a Perfect Independence Oracle. The inherent modularity of our approach paves the way for crucial future research, particularly focused on relaxing these assumptions to robustly handle unobserved confounders and work with finite, noisy real-world data.
Appendix A Additional Simulation Figures
Figure 3: Number of additional intervention rounds required when using method $r$ versus Adaptive_IP. Each panel corresponds to one of 9 networks; bars show the distribution of $r-\mathrm{IP}$ across seeds for different intervention limits $k_{\max}$.
Figure 4: Number of additional variables manipulations required when using method $r$ versus Adaptive_IP. Each panel corresponds to a graph-edge probability $p$; bars show the distribution of $r-\mathrm{IP}$ across seeds for different intervention limits $k_{\max}$ and network sizes $N$.
Figure 5: Number of additional variable manipulations required when using method $r$ versus Adaptive_IP. Each panel corresponds to one of 9 networks; bars show the distribution of $r-\mathrm{IP}$ across seeds for different intervention limits $k_{\max}$.
Appendix B Proofs & Theoretical Analysis
@@
In this appendix, we provide theoretical guarantees for the proposed adaptive intervention design algorithm (Algorithm 1). These results hold under the standard assumptions commonly made in causal discovery from interventions:
@@
Appendix C IP Model Extensions
C.1 Non-linear Cost Structures
Example (Two Variables):
Let $Y_{12}$ be a binary variable indicating whether both $V_{1}$ and $V_{2}$ are intervened on simultaneously ($Y_{12}=1$) or not ($Y_{12}=0$). Define a ”delta cost,” $C_{12}$, representing the extra cost incurred when both variables are jointly intervened on beyond their individual sums. In this example, $C_{12}=CI_{1,2}-CI_{1}-CI_{2}=10-1-1=8$.
The budget constraint becomes:
$CI_{1}\cdot X_{1}+CI_{2}\cdot X_{2}+C_{12}\cdot Y_{12}$ $\leq B$ (12)
$Y_{12}$ $\leq X_{1}$ (13)
$Y_{12}$ $\leq X_{2}$ (14)
$Y_{12}$ $\geq X_{1}+X_{2}-1$ (15)
These constraints ensure $Y_{12}=1$ if and only if both $X_{1}=1$ and $X_{2}=1$.
Generalization to Multiple Variables:
$Y_{S}$ $\leq X_{i_{j}},\quad\forall j$ (16)
$Y_{S}$ $\geq\sum_{j=1}^{k}X_{i_{j}}-(k-1)$ (17)
The budget constraint then incorporates these interactions:
$\sum_{i}CI_{i}\cdot X_{i}+\sum_{S}C_{S}\cdot Y_{S}$ $\leq B$ (18)
Physical Limitations and Prohibitive Costs:
C.2 Alternative Objective Functions
1. Average Gain over MEC:
2. Worst-Case Gain (Minimax Objective):
$\text{maximize}_{I}\quad\min_{G_{i}\in\text{MEC}(G^{*})}|R(I,G_{i})|$
3. Weighted Objective (Prioritizing Informative Edges):
How weights are assigned is crucial and heuristic. Examples include:
4. Targeted Discovery Objective:
5. Cost-Effectiveness (Approximation):
[…]
Here, $\lambda\geq 0$ is a penalty parameter balancing knowledge gain (potentially weighted) against intervention cost. This encourages finding solutions that are not just informative but also relatively cheap, though it requires tuning $\lambda$. Note that the budget constraint (Eq. 2) still applies.
C.3 Batch Interventions for Parallel Execution
The objective would typically aim to maximize the total knowledge gained across the batch:
Constraints need modification:
  • Budget per Experiment: A budget $B_{b}$ could apply to each experiment $b$.
    $\sum_{i\in V}(CI_{i}X_{i,b}+CO_{i}(1-X_{i,b}))\leq B_{b}\quad\forall b$
  • Intervention Limit per Experiment:
    $\sum_{i\in\mathcal{X}}X_{i,b}\leq k_{\text{max}}\quad\forall b$
    $\sum_{b=1}^{k_{\text{batch}}}\sum_{i\in V}CI_{i}X_{i,b}\leq B_{\text{total}}$
    $IDU_{ij,b}\leq O_{ij,b}+O_{ji,b}+A_{ij,b},\quad\dots\quad\forall b$
C.4 Dynamic Constraints